# FineSurE: Fine-grained Summarization Evaluation using LLMs

### v1

@@ Abstract
Automated evaluation is crucial for streamlining text summarization benchmarking and model development, given the costly and time-consuming nature of human evaluation. Traditional methods like ROUGE do not correlate well with human judgment, while recently proposed LLM-based metrics provide only summary-level assessment using Likert-scale scores. This limits deeper model analysis, e.g., we can only assign one hallucination score at the summary level, while at the sentence level, we can count sentences containing hallucinations. To remedy those limitations, we propose FineSurE, a fine-grained evaluator specifically tailored for the summarization task using large language models (LLMs). It also employs completeness and conciseness criteria, in addition to faithfulness, enabling multi-dimensional assessment. We compare various open-source and proprietary LLMs as backbones for FineSurE. In addition, we conduct extensive benchmarking of FineSurE against SOTA methods including NLI-, QA-, and LLM-based methods, showing improved performance especially on the completeness and conciseness dimensions. The code is available at https://github.com/DISL-Lab/FineSurE-ACL24.
FineSurE: Fine-grained Summarization Evaluation using LLMs
@@ 2 Related Work
With the emergence of LLMs, there is a move toward utilizing them as reference-free automated evaluators in diverse scenarios Shi et al. (2023); Lin and Chen (2023); Chen et al. (2023); Fu et al. (2023). Recently, a few efforts have been made to evaluate faithfulness using edited text Laban et al. (2023), atomic facts Min et al. (2023), and external knowledge base Feng et al. (2023), as well as to assess multi-dimensional aspects Liu et al. (2023); Shen et al. (2023); Wang et al. (2023). Although LLMs have shown promise as evaluators, they currently lack a fine-grained assessment and primarily focus on addressing faithfulness, without considering other important dimensions for high-quality summaries.
@@ 4.1.1 Faithfulness
Table 5: Performance of faithfulness evaluation using five open-source and four proprietary LLMs. The rightmost column is the success ratio of accurately following the prompt.
@@ 4.1.2 Completeness and Conciseness
Table 6: Performance of completeness and conciseness evaluation using five open-source and four proprietary LLMs. The rightmost column is the success ratio of accurately following the prompt.
@@
Table 7: Accuracy analysis of factuality error localization in assessing faithfulness using five open-source and four proprietary LLMs, where "Baseline" is the performance of random guessing. Top-1 values are marked in bold.
@@ Appendix A Factuality Error Type
Following the work Fu and Frank (2023), we use seven categories to define factuality error types, namely "out of context", "predicate," "entity," "circumstance," "coreference," "discourse link," and "grammatical". Table 8 provides the detailed description and example of the error categories.

### v2

# Fine-grained, Multi-dimensional Summarization Evaluation with LLMs

@@ Abstract
Automated evaluation plays a pivotal role in enhancing text summarization benchmarking and model development, given the high costs and time demands of human evaluation. Existing automated metrics like ROUGE restricts fine-grained and multi-dimensional analysis, such as identifying hallucinations at the sentence level or checking alignment at the keyfact level. To overcome these challenges, we introduce FineSurE, a fine-grained evaluation method tailored for summarization tasks using large language models (LLMs). FineSurE assesses summaries by using two distinct tasks, namely fact checking for faithfulenss evaluation and keyfact alignment for completeness and conciseness evaluation. We evalaute the performance of FineSurE by using various open-source and proprietary LLMs as the foundation. In addition, we benchmark its performance against state-of-the-art automated evaluators, including NLI-, QA-, and LLM-based approaches. Our extensive evaluations show that FineSurE exhibits higher agreements with human judgements for the three evaluation dimensions. The code is available at https://github.com/DISL-Lab/FineSurE-ACL24.
Fine-grained, Multi-dimensional Summarization Evaluation with LLMs
@@ 2 Related Work
With the emergence of LLMs, there is a move toward utilizing them as reference-free automated evaluators in diverse scenarios Shi et al. (2023); Lin and Chen (2023); Chen et al. (2023); Fu et al. (2023). Recently, a few efforts have been made to evaluate faithfulness using edited text Laban et al. (2023), atomic facts Min et al. (2023), and external knowledge base Feng et al. (2023), as well as to assess multi-dimensional aspects Liu et al. (2023); Shen et al. (2023); Wang et al. (2023). TrueTeacher Gekhman et al. employs LLMs to generate large-scale synthetic data for fact verification. Although LLMs have shown promise as evaluators, they currently lack a fine-grained assessment and primarily focus on addressing faithfulness, without considering other important dimensions for high-quality summaries.
@@ 4.1.1 Faithfulness
Table 5: Performance of faithfulness evaluation using six open-source and four proprietary LLMs. The rightmost column is the success ratio of accurately following the prompt.
@@ 4.1.2 Completeness and Conciseness
Table 6: Performance of completeness and conciseness evaluation using six open-source and four proprietary LLMs. The rightmost column is the success ratio of accurately following the prompt.
@@
Table 7: Accuracy analysis of factuality error localization in assessing faithfulness using six open-source and four proprietary LLMs, where "Baseline" is the performance of random guessing. Top-1 values are marked in bold.
@@ Appendix A Factuality Error Type
Following the work fu-frank-2023-seti, we use seven categories to define factuality error types, namely "out of context", "predicate," "entity," "circumstance," "coreference," "discourse link," and "grammatical". Table 8 provides the detailed description and example of the error categories.